IOPS requirement for RAID

An application has 1,000 heavy users at a peak of 2 IOPS each and 2,000
typical users at a peak of 1 IOPS each, with a read/write ratio of 2:1. It is
estimated that the application also experiences an overhead of 20 percent
for other workloads. Calculate the IOPS requirement for RAID 1, RAID 3,
RAID 5, and RAID 6.
Answer #1
not an answer but another questio
answered by: Timmy
Answer #2
not an answer but another question why did it become 90% ?
A general read/write ratio of 2:1
Maximum concurrency of active users at 90%


answered by: Rimmy
Answer #3
Dear,

1,000 heavy users at a peak of 2 IOPS each = 2000 IOPS

2,000 typical users at a peak of 1 IOPS each = 2000 IOPS

A general read/write ratio of 2:1

Maximum concurrency of active users at 90%

The requirement of [(2000+2000)* 0.9] = 3600 host-based IOPs during their peak activity period.

With 3600 total base IOPS and a 2:1 ratio:

For RAID 1/0:

(3600*2/3)+ (3600*1/3*2) =2400+2400 = 4800IOPS

For RAID 3:

(3600*2/3)+ (3600*1/3*0) =2400+0 = 2400IOPS

For RAID 5:

(3600*2/3)+ (3600*1/3*4) =2400+4800 = 7200IOPS

For RAID 6:

(3600*2/3)+ (3600*1/3*6) =2400+7200 = 9600IOPS

The application also experiences an overhead of 20 percent for other workloads.

RAID 1/0:

4800+20% = 4800+960 =5760IOPS

RAID3:

2400+20% = 2400+480 = 2880IOPS

RAID 5:

7200+20% = 7200 +1440 =8640IOPS

RAID 6:

9600+20% = 9600+1920 =11520IOPS

Hope this will helps you

answered by: Shuhan mirza